OpenStudy (jayden):

Find the coordinates of the midpoint of AB. A (12, -7) and B (-4, 7).

OpenStudy (anonymous):

Do you know the midpoint formula?

OpenStudy (jayden):

Yes xm=x1=x2/2 and ym=y1=y2/2

OpenStudy (jayden):

ow ok i get it thankyou

OpenStudy (jayden):

is the midpoint (4,0)

OpenStudy (anonymous):

Not quite. The formula for midpoint is; \[Midpoint=\left( \frac{ x_1+x_2 }{ 2 },\frac{ y_1+y_2 }{ 2 } \right)\]

OpenStudy (jayden):

(4,0) the answer

OpenStudy (anonymous):

Nope. Plug (12, -7) and (-4, 7) into the formula; \[Midpoint=\left( \frac{ 12-4 }{ 2 },\frac{ -7+7 }{ 2 } \right)\] Try now..

OpenStudy (anonymous):

Oh, wow big mistake on my part, haha. Yes you are right.

OpenStudy (jayden):

ok